# Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation

### v1

@@ Abstract
Variational Autoencoders (VAEs) with global priors mirror the training set’s class frequency in latent space, underrepresenting tail classes and reducing generative fairness on imbalanced datasets. While $t^{3}$VAE improves robustness via heavy-tailed Student’s t-distribution priors, it still allocates latent volume proportionally to the class frequency. In this work, we address this issue by explicitly enforcing equitable latent space allocation across classes. To this end, we propose Conditional-$t^{3}$VAE, which defines a per-class Student’s t joint prior over latent and output variables, preventing dominance by majority classes. Our model is optimized using a closed-form objective derived from the $\gamma$-power divergence. Moreover, for class-balanced generation, we derive an equal-weight latent mixture of Student’s t-distributions. On SVHN-LT, CIFAR100-LT, and CelebA, Conditional-$t^{3}$VAE consistently achieves lower FID scores than both $t^{3}$VAE and Gaussian-based VAE baselines, particularly under severe class imbalance. In per-class F1 evaluations, Conditional-$t^{3}$VAE also outperforms the conditional Gaussian VAE across all highly imbalanced settings. While Gaussian-based models remain competitive under mild imbalance ratio ($\rho\lesssim 3$), our approach substantially improves generative fairness and diversity in more extreme regimes.
@@ 1 Introduction
Class imbalance and long-tail distributions are prevalent in real-world datasets, yet generative models often fail to represent rare classes accurately. When trained on skewed data, these models tend to overfit dominant modes and underrepresent minority ones in latent and output spaces, resulting in biased or unfair generations. This issue is especially critical in sensitive applications such as facial synthesis [1] and medical imaging [2], where such biases can exacerbate social and diagnostic disparities [3].
Variational Autoencoders (VAEs) [4] are a widely used class of generative models, valued for their probabilistic formulation, stable training, and compatibility with latent-variable modeling frameworks leading in image quality such as Latent Diffusion Models (LDMs) [5]. Although GANs and diffusion models often achieve lower FID scores, VAEs offer unique advantages in class-conditional generation, interpretability, and efficient inference, making them strong candidates for improving fairness under class imbalance. Standard VAEs commonly use isotropic Gaussian priors, which inadequately model heavy-tailed structures and rare phenomena [6]. Prior efforts to address this have introduced non-Gaussian priors, particularly Student’s t-distributions [7, 8, 9, 10], to enhance robustness. However, these approaches often rely on global priors, causing the latent space to be dominated by majority classes under skewed distributions.
We address this issue with Conditional-$t^{3}$VAE (C-$t^{3}$VAE), a conditional generative model that imposes a per-class Student’s t-distribution prior over the joint latent-output space. This design ensures allocating an equal latent space volume per class, thereby mitigating majority class dominance, while the heavy tails of the Student’s t-distribution more effectively capture intra-class variation. To enable class-balanced sampling, we introduce an equal-weight mixture of Student’s t-distributions with analytically derived component variances. Together, these components enable balanced class-conditional generation and mitigate bias present in unconditional models. We summarize our main theoretical and empirical contributions in the following points :
@@
  • We outperform relevant baselines in FID on SVHN-LT [11], CIFAR100-LT [12], and CelebA [13] under severe imbalance, and show via per-class evaluation that C-$t^{3}$VAE better avoids mode collapse, exceeding a conditional VAE in per class Recall and F1 while remaining competitive on Precision.
  • We identify the imbalance ratio threshold $\rho\approx 3$, beyond which Gaussian priors become suboptimal, providing guidance for model selection on skewed datasets.
2 Related Work
Since the introduction of VAEs [4], many extensions have sought to improve latent representation by replacing the standard Gaussian prior with more expressive alternatives. These include Gaussian mixtures [14, 15], hyperspherical priors [16], normalizing flows [17], Riemannian priors [18], and implicit distributions [19]. While most retain the Evidence Lower Bound (ELBO) formulation, others adopt alternative objectives or divergence measures for added flexibility.
To address long-tailed or imbalanced data, Student’s t-distributions have been explored for their robustness and heavy tails [6]. Methods such as [7] and [8] model the latent space of the autoencoder through a t-distributed prior, and rely on KL-divergence-based ELBO objectives. However, since the KL divergence lacks a closed-form solution for the Student’s t-distributions, these methods resort to numerical approximation. The $t^{3}$VAE [10] improves on this by modeling the joint latent-output distribution and optimizing a closed-form $\gamma$-divergence objective [9]. Nonetheless, it still employs a global latent prior, which results in a latent space volume allocation reflecting class frequency and leading to imbalance present in the generated samples.
Other works address fairness in generative modeling using normalizing flows [17] and diffusion models [20, 21]. However, since leading image generation models are based on Latent Diffusion Models [5] and VAEs are a corner-stone in the design of these models, we deem it crucial to improve VAEs’ ability to handle imbalanced data. Consequently, as fairness in latent space allocation remains under-explored in VAEs, we introduce a class-conditional, heavy-tailed prior to address this gap and allow for balanced latent space allocation across classes.
@@ 3 Background
This section introduces the theoretical background and baseline models relevant to our work. We assume access to a labeled, imbalanced dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{n}$ is a data sample of dimension $n$, $y_{i}\in\{1,\dots,K\}$ its class label and $m$ the latent space dimension.
3.1 VAEs and Conditional-VAEs
VAEs [4] are generative models trained via variational inference by maximizing the ELBO of the log-likelihood. The standard objective of this model is
@@ 3.1 VAEs and Conditional-VAEs
where the first term is the reconstruction loss with $p_{\theta}(x|z)$ being the decoder. The second term is the Kullback–Leibler (KL) divergence between the approximate posterior $q_{\phi}(z|x)$ and the prior $p(z)$. The $\beta$-VAE is a weighted variant of the VAE model which introduces a $\beta$ scaling term for the KL divergence [22]:
@@
with $p(z)\sim\mathcal{N}_{m}(0,I)$, $q_{\phi}(\cdot|x)\sim\mathcal{N}_{m}(\mu_{\phi}(x),\Sigma_{\phi}(x))$, and $p_{\theta}(x|z)\sim\mathcal{N}_{m}(\mu_{\theta}(z),\sigma^{2}I)$. $\mu_{\phi}(\cdot)$ and $\Sigma_{\phi}(\cdot)$ are the mean and the covariance matrices inferred through a neural network with parameter $\phi$ given the input $x$. This variant of the VAE model allows to place more weight on disentangling the latent space or on the reconstruction of the data points. To generate samples from the VAE or the $\beta$-VAE model, we sample a latent vector $z\sim\mathcal{N}_{m}(0,I)$. Then, the generated data point would be $\hat{x}\sim\mathcal{N}_{m}(\mu_{\theta}(z),\sigma^{2}I)$.
Nevertheless, since Eq. (1) and Eq. (2) optimize the ELBO over the data distribution $p_{data}(x)$, which can be decomposed as $p_{\mathrm{data}}(x)=\sum_{y_{i}}p(y_{i})\,p_{\mathrm{data}}(x\mid y_{i}),$ this optimization inherently biases the model toward head classes with larger $p(y_{i})$. As a result, most generated samples come from overrepresented classes, while tail classes’ samples are underrepresented and of lower quality, a phenomenon commonly referred to as mode collapse. Therefore, when labels are available, it is preferable to define class-conditional approximate posterior and prior distributions: $q_{\phi}(z|x,y)$ and $p(z|y)$. This yields the Conditional-VAE (CVAE) model trained using the objective [23]:
@@
Here, in Eq. (3) we constrain all $p(y_{i})$ to be equal and omit them from the loss function in order not to exacerbate the issue of class imbalance in the latent space. Also, we define $p(z|y)\sim\mathcal{N}_{m}(\mu_{y},I)$ with learnable class-wise means $\mu_{y}$. To generate a data point $\hat{x}_{y}$ from class $y$, we sample $z_{y}\sim\mathcal{N}_{m}(\mu_{y},I)$, then we get $\hat{x}_{y}\sim p_{\theta}(x|z_{y},y)$. Nevertheless, despite conditioning, this formulation remains Gaussian. Unlike Student’s t-distributions, Gaussian priors poorly approximate heavy-tailed data distributions [6].
3.2 Multivariate Student’s t-Distribution and $\gamma$-Power Divergence
@@
The power form of this distribution prevents a closed-form KL divergence between two Student’s t-distributions. Instead, the $\gamma$-power divergence $\mathcal{D}_{\gamma}(q\|p)$ is used [9, 10]. This divergence is defined for $q\sim t_{d}(\mu_{0};\Sigma_{0};\nu)$, $p\sim t_{d}(\mu_{1};\Sigma_{1};\nu)$ starting from the $\gamma$-entropy $\mathcal{H}_{\gamma}(p)$ and cross-entropy $\mathcal{C}_{\gamma}(q,p)$
@@
$\mathcal{D}_{\gamma}(q\|p):=\gamma^{-1}\left(\mathcal{C}_{\gamma}(q,p)-\mathcal{H}_{\gamma}(p)\right)$ (5)
with $\gamma=-\frac{2}{\nu+d}$. Then, substituting the definition of a Student’s t-distribution from Eq. (4) into Eq. (5), the following closed-form formula for the $\gamma$-power divergence can be derived (See Appendix A):
@@
$\quad+|\Sigma_{1}|^{-\frac{\gamma}{2}}|\Sigma_{0}|^{\frac{\gamma^{2}}{2(1+\gamma)}}\Bigg{(}1+\frac{\operatorname{Tr}\left(\Sigma_{1}^{-1}\Sigma_{0}\right)}{\nu-2}+\frac{(\mu_{0}-\mu_{1})^{\top}\Sigma_{1}^{-1}(\mu_{0}-\mu_{1})}{\nu}\Bigg{)}\Bigg{]}.$ (6)
@@ 3.3.1 Definition
The $t^{3}$VAE model [10] is a non-ELBO-based autoencoder which models the joint prior distribution $p_{\theta}(x,z)$ using multivariate Student’s t-distributions
$p_{\theta}(x,z)\propto\sigma^{-n}\left[1+\frac{1}{\nu}\left(\|z\|^{2}+\frac{\|x-\mu_{\theta}(z)\|^{2}}{\sigma^{2}}\right)\right]^{-\frac{\nu+m+n}{2}},$
where $\mu_{\theta}(\cdot)$ is the decoder neural network with parameter $\theta$ whereas $\sigma$ is a parameter controlling the decoder’s output covariance. From this joint distribution, the marginal latent prior $p(z)$ and decoder distribution $p_{\theta}(x|z)$ can be defined. Furthermore, the approximate posterior distribution is defined as :
@@
with $\gamma=-\frac{2}{\nu+n+m}$ and $C_{1}$ and $C_{2}$ being theoretically derived constants. We note that the first term in this loss function represents the standard reconstruction term in VAE models and the rest of the terms are regularization terms over the latent space. To sample from the latent space of the $t^{3}$VAE, [10] propose the $p_{\nu}^{\star}(z)=t_{m}(0,\tau^{2}I,\nu+n)$ distribution with
$\tau^{2}=\frac{1}{1+\nu^{-1}n}\left(\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{\frac{2}{\nu+n-2}}.$ (8)

### v2

# Heavy-Tailed Class-Conditional Priors for Long-Tailed Generative Modeling

@@ Abstract
Variational Autoencoders (VAEs) with global priors trained under an imbalanced empirical class distribution can lead to underrepresentation of tail classes in the latent space. While $t^{3}$VAE improves robustness via heavy-tailed Student’s $t$-distribution priors, its single global prior still allocates mass proportionally to class frequency. We address this latent geometric bias by introducing C-$t^{3}$VAE, which assigns a per-class Student’s $t$ joint prior over latent and output variables. This design promotes uniform prior mass across class-conditioned components. To optimize our model we derive a closed-form objective from the $\gamma$-power divergence, and we introduce an equal-weight latent mixture for class-balanced generation. On SVHN-LT, CIFAR100-LT, and CelebA datasets, C-$t^{3}$VAE consistently attains lower FID scores than $t^{3}$VAE and Gaussian-based VAE baselines under severe class imbalance while remaining competitive in balanced or mildly imbalanced settings. In per-class F1 evaluations, our model outperforms the conditional Gaussian VAE across highly imbalanced settings. Moreover, we identify the mild imbalance threshold $\rho<5$, for which Gaussian-based models remain competitive. However, for $\rho\geq 5$ our approach yields improved class-balanced generation and mode coverage.
@@ 1 Introduction
Class imbalance and long-tail distributions are common in real-world datasets, yet generative models often fail to represent rare classes accurately. Under skewed training data, models tend to overfit dominant modes and underrepresent minority ones in latent and output spaces, resulting in biased generations. This issue is particularly consequential in applications such as facial synthesis (Mehta et al., 2024) and medical imaging (Pinaya et al., 2022), where such biases can exacerbate social and diagnostic disparities (Naik & Nushi, 2023). Addressing class imbalance to ensure balanced representational capacity across categories remains a major challenge for all generative models.
For Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), their inherently unstable training dynamics, exacerbated by data imbalance, often lead to biased and mode-collapsed generations. Several works have sought to mitigate these effects. The Wasserstein GAN (Arjovsky et al., 2017) replaces the Jensen–Shannon divergence with the Wasserstein distance to stabilize optimization and improve sample diversity. PacGAN (Lin et al., 2018) enhances robustness by packing multiple samples into the discriminator, reducing mode collapse and improving diversity under implicit imbalances. Similarly, (Asokan & Seelamantula, 2020) introduces negative data augmentation to prevent under-representation of minority classes, enabling more class-balanced supervised GAN training. More recently, RareGAN (Lin et al., 2022) addresses unlabeled long-tailed data through a weighted loss and adaptive labeling budgets, improving both balance and generative diversity. Despite these advances, GANs remain difficult to train reliably, especially under strong imbalance.
Diffusion models have emerged as a more stable alternative to GANs, achieving superior image quality and convergence behavior (Dhariwal & Nichol, 2021). Several diffusion-based methods explicitly tackle data imbalance. Class-Balancing Diffusion Models (Qin et al., 2023) modify the denoising process to be class-invariant, while (Zhang et al., 2024) employs weighted score matching with Bayesian calibration to transfer knowledge from majority to minority classes, improving diversity and fidelity on long-tailed datasets. Heavy-Tailed Diffusion Models (Pandey et al., 2025) take a different approach by replacing the Gaussian noise assumption with a Student’s $t$ formulation, offering a more robust fit for imbalanced distributions. Nevertheless, effectively addressing class imbalance within diffusion frameworks remains an open and active research area.
Variational Autoencoders (VAEs) (Kingma & Welling, 2013) are valuable models when interpretable, structured latent spaces and efficient single-step sampling are needed. VAEs provide a principled probabilistic foundation, stable training dynamics, and compatibility with latent-variable pipelines such as latent diffusion (Rombach et al., 2022). However, in class imbalanced setting, standard VAEs using isotropic Gaussian priors struggle to capture heavy-tailed or rare structures, creating class-coverage bottlenecks (Tam & Dunson, 2025). Therefore, prior work explores non-Gaussian priors, especially Student’s $t$-distributions (Takahashi et al., 2018; Abiri & Ohlsson, 2020; Eguchi, 2021; Kim et al., 2024), to improve robustness to outliers and class imbalance.
In this work, we view class imbalance not only as a sampling issue but as a geometric one: optimizing under the empirical distribution tends to allocate latent probability mass in proportion to class frequency. Consequently, majority classes occupy larger effective regions of latent space, while minority classes are confined to smaller regions with reduced representational capacity. This interpretation is particularly transparent in latent-variable models with an explicit global prior, such as VAEs, where representational allocation is directly governed by prior density. While imbalance also affects GANs and diffusion models, in those frameworks it is intertwined with adversarial training dynamics or iterative score-based denoising. In contrast, VAEs allow us to study how prior specification alone shapes latent geometry. We therefore focus on VAEs to isolate the role of prior-induced frequency bias in a controlled and analytically tractable setting.
Current VAE based approaches retain a single global prior (Kingma & Welling, 2013; Takahashi et al., 2018; Mathieu et al., 2019b; Abiri & Ohlsson, 2020). Hence through our geometric view the latent remains frequency-aligned where class regions scale with empirical probability and heavy tails alone do not eliminate this frequency-induced geometric bias. Indeed, optimizing a VAE under $p_{data}(x)=\sum_{y}p(y)p(x|y)$ tends to allocate latent regions whose effective volume correlates with class frequency $p(y)$. We address this issue by introducing the Conditional-$t^{3}$VAE (C-$t^{3}$VAE), which imposes a per-class Student’s $t$-distribution prior over the joint latent-output space. The design promotes uniform prior mass across class-conditioned components under the sampling distribution, mitigating majority-class dominance while the heavy tails capture intra-class variation. For class-balanced generation, we derive an equal-weight latent mixture of Student’s $t$-distributions with analytically derived component variances. Importantly, C-$t^{3}$VAE class-specific heavy-tailed priors induce a structured latent mixture whose geometry differs fundamentally from both global heavy-tailed priors and Gaussian conditional VAEs. We summarize our main contributions in the following points :
@@
  • We outperform relevant baselines in FID on SVHN-LT (Netzer et al., 2011), CIFAR100-LT (Cao et al., 2019), and CelebA (Liu et al., 2015) under severe imbalance, and show via per-class evaluation that C-$t^{3}$VAE better avoids mode collapse, exceeding a Gaussian conditional VAE in per class Recall and F1 while remaining competitive on Precision.
  • We empirically observe a transition regime around $\rho\approx 5$, beyond which Gaussian priors become suboptimal, providing guidance for model selection on skewed datasets.
2 Related works
Since the introduction of Variational Autoencoders (VAEs) (Kingma & Welling, 2013), numerous extensions have aimed to enhance latent representation expressiveness by replacing the standard Gaussian prior with more flexible formulations. Notable examples include Gaussian mixtures (Dilokthanakul et al., 2016; Saseendran et al., 2021), hyperspherical priors (Davidson et al., 2018), normalizing flows (Jaini et al., 2020), Riemannian manifolds (Chadebec et al., 2023), and implicit distributions (Takahashi et al., 2019). Most of these retain the Evidence Lower Bound (ELBO) optimization framework, while others introduce alternative divergence measures for greater modeling flexibility (Makhzani et al., 2016). Hierarchical VAEs (Vahdat & Kautz, 2020) and vector-quantized VAEs (van den Oord et al., 2017) have also been proposed to improve disentanglement and mitigate posterior collapse.
Student’s $t$-distributions have been widely studied for their heavy-tailed robustness to outliers (Tam & Dunson, 2025). Early works (Mathieu et al., 2019a; Abiri & Ohlsson, 2020) incorporated $t$-distributed priors into the VAE framework using KL-divergence-based ELBO objectives to promote robust latent encodings. However, because the KL divergence between $t$-distributions lacks a closed-form solution, such approaches depend on numerical approximations, increasing computational cost. (Takahashi et al., 2018) proposes a Student’s $t$ decoder to improve generative performance, but the heavy-tailed assumption is confined to the decoder, leaving the latent space Gaussian and limiting validation to tabular data. The $t^{3}$VAE (Kim et al., 2024) further advances this line of work by jointly modeling latent and output distributions under Student’s $t$-assumptions for the encoder, decoder, and prior, replacing the KL divergence with a closed-form $\gamma$-divergence (Eguchi, 2021).
While several works improve prior expressiveness or robustness, they do not explicitly analyze how global prior formulations induce frequency-aligned allocation of latent probability mass under imbalanced data. In particular, existing approaches focus on better density modeling or optimization stability, rather than controlling how representational capacity is distributed across classes. As a result, even expressive or heavy-tailed priors may continue to reflect empirical class proportions in latent space.
A complementary line of research addresses imbalance through conditional modeling. Conditional VAEs (CVAEs) (Kingma et al., 2014; Sohn et al., 2015) condition latent encodings on class labels, enabling targeted generation of minority categories. However, their Gaussian priors poorly approximate heavy-tailed or rare data structures. Our C-$t^{3}$VAE bridges two threads: it inherits the heavy-tailed robustness of $t^{3}$VAE while adopting the class-conditional structure of CVAEs, and further introduces a theoretically grounded equal-weight mixture sampling scheme that explicitly counteracts majority-class dominance at generation time. Hence, our contribution is not merely heavy tails nor conditioning, but decoupling representational capacity from empirical frequency through prior mass allocation.
@@ 3 Background
This section introduces the theoretical background and baseline models relevant to our work. We assume access to a labeled, imbalanced dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{n}$ is a data sample of dimension $n$, $y_{i}\in\{1,\dots,K\}$ its class label with $K$ being the number of classes and $m$ being the latent space dimension.
3.1 VAEs and Conditional VAEs
VAEs (Kingma & Welling, 2013) are generative models trained via variational inference by maximizing the Evidence Lower Bound (ELBO) of the log-likelihood. The standard objective of this model is
@@ 3.1 VAEs and Conditional VAEs
where the first term is the reconstruction loss with $p_{\theta}(x|z)$ being the decoder model. The second term is the Kullback–Leibler (KL) divergence between the approximate posterior $q_{\phi}(z|x)$ and the prior $p(z)$. The $\beta$-VAE is a weighted variant of the VAE model which introduces a $\beta$ scaling term for the KL divergence (Higgins et al., 2017):
@@
with $p(z)\sim\mathcal{N}_{m}(0,I)$, $q_{\phi}(\cdot|x)\sim\mathcal{N}_{m}(\mu_{\phi}(x),\Sigma_{\phi}(x))$, and $p_{\theta}(x|z)\sim\mathcal{N}_{m}(\mu_{\theta}(z),\sigma^{2}I)$. $\mu_{\phi}(\cdot)$ and $\Sigma_{\phi}(\cdot)$ are the mean and covariance matrices of the approximate posterior. They are inferred by a neural network with parameters $\phi$ given the input $x$. Moreover, $\mu_{\theta}(\cdot)$ is the decoder neural network with parameter $\theta$ and $\sigma$ is a parameter controlling the decoder’s output covariance. This variant of the VAE model allows to place more weight on disentangling the latent space or on the reconstruction of the data points. To generate samples from the VAE or the $\beta$-VAE model, we sample a latent vector $z\sim\mathcal{N}_{m}(0,I)$. Then, the generated data point would be $\hat{x}\sim\mathcal{N}_{m}(\mu_{\theta}(z),\sigma^{2}I)$.
Nevertheless, since Eq. (1) and Eq. (2) optimize the ELBO over the data distribution $p_{data}(x)$, which can be decomposed as $p_{\mathrm{data}}(x)=\sum_{y_{i}}p(y_{i})\,p_{\mathrm{data}}(x\mid y_{i})$. In the context of imbalanced data, this optimization inherently biases the model toward head classes with larger $p(y_{i})$. As a result, most generated samples come from overrepresented classes, while tail classes’ samples are underrepresented and of lower quality, a phenomenon commonly referred to as mode collapse. Therefore, when labels are available, it is preferable to define class-conditional posterior and prior distributions: $q_{\phi}(z|x,y)$ and $p(z|y)$. This yields the Conditional-VAE (CVAE) model trained using the objective (Kingma et al., 2014):
@@
In practice, in Eq. (3) we optimize the class-conditional objective with uniform weighting across labels, removing explicit frequency-dependent scaling from the loss. Also, we define $p(z|y)\sim\mathcal{N}_{m}(\mu_{y},I)$ with learnable class-wise means $\mu_{y}$. To generate a data point $\hat{x}_{y}$ from class $y$, we sample $z_{y}\sim\mathcal{N}_{m}(\mu_{y},I)$, then we get $\hat{x}_{y}\sim p_{\theta}(x|z_{y},y)$. Nevertheless, despite conditioning, this formulation remains Gaussian. Unlike Student’s $t$-distributions, Gaussian priors poorly approximate heavy-tailed data distributions (Tam & Dunson, 2025).
3.2 Multivariate Student’s $t$-Distribution
@@
The power form of this distribution prevents a closed-form KL divergence between two Student’s $t$-distributions. Instead, the $\gamma$-power divergence $\mathcal{D}_{\gamma}(q\|p)$ is used (Eguchi, 2021; Kim et al., 2024). This divergence is defined for $q\sim t_{d}(\mu_{0};\Sigma_{0};\nu)$, $p\sim t_{d}(\mu_{1};\Sigma_{1};\nu)$ as
$\mathcal{D}_{\gamma}(q\|p):=\frac{\mathcal{C}_{\gamma}(q,p)-\mathcal{H}_{\gamma}(p)}{\gamma}$ (5)
with $\gamma=-\frac{2}{\nu+d}$, the $\gamma$-entropy $\mathcal{H}_{\gamma}(p)$ and $\gamma$-cross-entropy $\mathcal{C}_{\gamma}(q,p)$ being
@@
Then, substituting the definition of a Student’s $t$-distribution from Eq. (4) into Eq. (5), the following closed-form formula for the $\gamma$-power divergence can be derived (Derivation of Proposition 3 in (Kim et al., 2024))
@@
$\quad+|\Sigma_{1}|^{-\frac{\gamma}{2(1+\gamma)}}\Bigg(1+\frac{\operatorname{Tr}\left(\Sigma_{1}^{-1}\Sigma_{0}\right)}{\nu-2}+\frac{(\mu_{0}-\mu_{1})^{\top}\Sigma_{1}^{-1}(\mu_{0}-\mu_{1})}{\nu}\Bigg)\Bigg].$ (6)
@@ 3.3.1 Definition
The $t^{3}$VAE model (Kim et al., 2024) is a non-ELBO-based autoencoder which models the joint prior distribution $p_{\theta}(x,z)$ using multivariate Student’s $t$-distribution
$p_{\theta}(x,z)=\sigma^{-n}C_{\nu,m+n}\left[1+\frac{1}{\nu}\left(\|z\|^{2}+\frac{\|x-\mu_{\theta}(z)\|^{2}}{\sigma^{2}}\right)\right]^{-\frac{\nu+m+n}{2}}.$
From this joint distribution, the marginal latent prior $p(z)$ and decoder distribution $p_{\theta}(x|z)$ can be defined as follows
$p(z)=t_{m}(z|0,I,\nu),\quad p_{\theta}(x|z)=t_{n}\left(x\middle|\mu_{\theta}(z),\frac{1+\nu^{-1}\|z\|^{2}}{1+\nu^{-1}m}\sigma^{2}I,\nu+m\right).$
Furthermore, the posterior distribution is defined as :
@@
with $\gamma=-\frac{2}{\nu+n+m}$ and $C_{1}$ and $C_{2}$ being constants theoretically derived in (Kim et al., 2024). We note that the first term in this loss function represents the standard reconstruction term in VAE models and the rest of the terms are regularization terms over the latent space. To sample from the latent space of the $t^{3}$VAE, (Kim et al., 2024) propose the $p_{\nu}^{\star}(z)=t_{m}(0,\tau^{2}I,\nu+n)$ distribution with
$\tau^{2}=\left(1+\nu^{-1}n\right)^{-1}\left(\sigma^{n}C^{-1}_{\nu,n}\quad\frac{\nu+n-2}{\nu-2}\right)^{-\frac{2}{\nu+n-2}}.$ (8)